Draw the balanced reaction of salicylic acid with 2 moles of NaOH. Don't forget counterions and formal charges.
Answer #1

The structure of salicylic acid is as follows:

image.png

O=C(O)c1ccccc1O

salicylic acid

Salicylic acid has a carboxylic acid (-COOH) group that readily reacts with the base and forms its sodium salt. Salicylic acid also has a phenolic -OH that acts as a weak acid and reacts with the base and forms its respective sodium salt. Finally, a disodium salt is formed.

The reaction of salicylic acid with excess of \(\mathrm{NaOH}\) is as follows:

image.png

O=C(O)c1ccccc1O

— 2NaOH  —>

O=C(O[Na])c1ccccc1[N+](=O)[O-]

answered by: Thrilled